# SLIFER: Investigating Performance and Robustness of Malware Detection Pipelines

### v1

@@ 5.2. SLIFER Classification
Table 4. True Positive Rate (TPR) of SLIFER divided by malware family inside in $\mathcal{D}_{1}$
@@
Table 5. True Positive Rate (TPR) of SLIFER divided by malware family inside in $\mathcal{D}_{2}$
@@
As for differences between the pipeline with the dynamic module and without, we can see a negligible improvement in TPR and only in with $\mathcal{D}_{2}$ at the expense of FPR for both datasets. Also, to better understand such an improvement, we display the distinct predictive capabilities of both strategies on the different families that compose our datasets, by showing results in Table 4 for $\mathcal{D}_{1}$ and Table 5 for $\mathcal{D}_{2}$. For the first dataset, which is severely imbalanced in terms of malware families, we can notice an identical detection rate when treating errors as benignware. Also, the number of errors is very low, with a peak of 3% (40 samples) on Grayware programs, there is no clear advantage brought by dynamic analysis on this dataset.
@@
Similarly, for the second dataset, both techniques score identical results, with low TPR for Dropper and Keylogger families. On this dataset, we confirm what has already been stated by Dambra et al.(Dambra et al., 2023), that dynamic analysis marginally helps in detecting specific families that are harder to detect by static analysis. However, by also looking at numbers, all the emulation errors are also focused on those families as well. Thus, it is possible that these numbers would further improve with a better emulation or virtualization tool as back-end for the dynamic analysis module. We derive that adding a dynamic module at the end of a sequential pipeline for malware detection might not improve performances as much as expected. In particular, the inclusion of this module might increase the errors encountered at analysis time, which must be dealt with accordingly. Secondly, differently from Dambra et al. which consider models with static and dynamic features concatenated together, dynamic analysis as a separate part of a detection pipeline does not increase the capabilities of the overall system. Hence, we can answer RQ2, by stating the sequential architecture of SLIFER improves TPR within a low FPR regime than single models alone, but not really thanks to the dynamic module.
@@
Take-home message 2: SLIFER outperforms single models in terms of TPR, by keeping low FPR as well. SLIFER also reduces the number of pre-processing errors thanks to detection achieved by earlier modules. However, we complement the findings posed by Dambra et al. (Dambra et al., 2023) by highlighting the minimal increment in performance brought by dynamic analysis in binary classification.
5.3. Robustness Against Transfer Attacks
@@
We now answer RQ3 by testing the robustness of our architecture against adversarial EXEmples computed on GBDT-EMBER and transferred again the SLIFER, and we show results in Table 6. From the randomly-sampled 1k malicious samples, GAMMA is able to produce 388 EXEmples that bypass GBDT-EMBER detection. We use this set to compute the robustness against transfer attacks of SLIFER with and without the dynamic module. Since we are interested also in the effect that signatures play in countering malware, we want to quantify how much they can be useful when dealing with adversarial EXEmples. Thus, we also add to the comparison a version of SLIFER where YARA pattern-matching is removed (SLIFER No Sign.). Lastly, to understand whether an attack that manipulate only the structure of a program might also affect a behavioral model, we also test the robustness of Nebula alone.
@@
Table 6. Adversarial Detection Rate (ADR) of diverse architectures/model under adversarial attacks. We compare not manipulated malware samples and the same samples after GAMMA manipulation (Adversarial EXEmples).
While all untainted malware samples are detected by SLIFER, with the exception of Nebula alone due to the high number of emulation errors, results of transfer evaluation highlight an evasive trend against all models. These suggest that GAMMA attack can evade the whole SLIFER, decreasing the Detection Rate by 14%. This is achieved also thanks to the policy we established of flagging samples as benignware in case of errors of the last dynamic module. As we can see from the ADR if we count errors as malware, we derive that the 7% of Adversarial EXEmples crashes the analysis, and the remaining 7% bypasses SLIFER without pre-processing errors. By looking at Nebula performances on these data, it is likely that roughly half of them are not detected at all, but also that some are, surprisingly, evading detection. This can be noticed by looking again at the results scored by Nebula: when considering errors as malware, there is a small drop of performance (3 %) implying that those EXEmples do not crash the pipeline. By inspecting the results, we discovered that (i) some malware samples obtain a different score when emulated multiple times inside a single Speakeasy session, and (ii) GAMMA changes the output of Nebula. We analysed those reports, and we discovered that samples that invoke specific Windows APIs like:
@@
  • GetModuleFileNameW retrieves a different file name. It is likely that Speakeasy rename the file to analyse with its hash. Thus, the output report of the untainted malware and its adversarial EXEmple counterpart are different.
This analysis is not comprehensive, as it would need deeper reverse engineering of those samples, but, even if for few samples, we can conclude that also dynamic analysis is evaded by non-behavioral attacks, not strictly due to the effect of perturbations. Lastly, 5 samples that originally caused errors during emulation are now analysed by Speakeasy after the manipulation. This result needs further investigation, but they bring evidence that dynamic analysis based on emulation can be weak and discontinuous not only against adversarial transfer attacks, but also to multiple evaluations of the same samples.
If we remove dynamic analysis from SLIFER (SLIFER No Dyn. in), we detect a drop of ADR by 10%, implying that such a module was indeed able to stop most of the adversarial EXEmples. However, counter-intuitively, if we instead remove signatures (SLIFER No Sign.) and we keep dynamic analysis, we assist to a drop of ADR by 15% w.r.t. SLIFER. This means that, surprisingly, pattern-matching with YARA rules has a bigger impact on robustness than dynamic analysis itself. This relevant contribution of YARA rules is due to how the section injection manipulation works: while mixing together sections extracted from input data, it is likely that GAMMA creates some patterns that trigger YARA rules, contrary to the original PE. We notice that 20 samples evade the YARA module before GAMMA manipulation, but they are detected once perturbed. To better understand this result, we analyse a small subset of rules that are triggered by adversarial EXEmples, and we analyse them through expert domain knowledge. Through this study, we isolate culprits of such triggers:
@@
Thus, we can conclude that SLIFER is indeed robust against transfer adversarial EXEmples computed against its strongest component, but such robustness is achieved more by static signatures than dynamic analysis alone.
@@
Take-home message 3: SLIFER can withstand adversarial transfer attacks, but the effect of dynamic analysis is only and counter-intuitively marginal. Attacks can trigger YARA rules that were not predictive against the original untainted samples, since GAMMA introduces artifacts.
5.4. Computation Times Comparison

### v3

@@ 5.2 SLIFER Classification Performance
Table 5: True Positive Rate (TPR) of non-calibrated SLIFER divided by malware family inside in $\mathcal{D}_{2}\,$. Results in brackets are calculated considering detection errors as malware. They are considered benignware otherwise.
@@
As for differences between the pipeline with the dynamic module and without, we can see a negligible improvement in TPR and only in with $\mathcal{D}_{2}\,$ at the expense of FPR for both datasets.
@@
Also, to better understand such an improvement, we display the distinct predictive capabilities of both strategies on the different families that compose our datasets, by showing results in Table 4 for $\mathcal{D}_{1}\,$ and Table 5 for $\mathcal{D}_{2}\,$. For the first dataset, which is severely imbalanced in terms of malware families, we can notice an identical detection rate when treating errors as benignware. Also, the number of errors is very low, with a peak of 3% (40 samples) on Grayware programs, there is no clear advantage brought by dynamic analysis on this dataset.
@@
Similarly, for the second dataset, both techniques score identical results, with low TPR for Dropper and Keylogger families. Furthermore, this trend is confirmed by the F1-Score performed by the pipeline. On this dataset, we confirm what has already been stated by (Dambra et al., 2023), that dynamic analysis marginally helps in detecting specific families that are harder to detect by static analysis. However, by also looking at numbers, all the emulation errors are also focused on those families as well. Thus, it is possible that these numbers would further improve with a better emulation or virtualization tool as back-end for the dynamic analysis module. We derive that adding a dynamic module at the end of a sequential pipeline for malware detection might not improve performances as much as expected. In particular, this module might increase the errors encountered at analysis time, which must be dealt with accordingly. Secondly, different from Dambra et al. which consider models with static and dynamic features concatenated together, dynamic analysis as a separate part of a detection pipeline does not increase the capabilities of the overall system.
Hence, we can answer RQ2, by stating the sequential architecture of SLIFER improves TPR within a low FPR regime than single models alone, but not thanks to the dynamic module.
@@
Take-home message 2: SLIFER outperforms single models in terms of the considered metrics, by also reducing the number of pre-processing errors. We complement previous findings (Dambra et al., 2023) by highlighting the minimal increment in performance brought by dynamic analysis in binary classification.
5.3 SLIFER Performance After Calibration
Here we present the calibration process of SLIFER, and how it changes the performance of all the considered models.
Performance on Validation Set. As described in Sect. 4, we perform an exhaustive grid search to select the best combination of decision thresholds characterizing SLIFER’s models. We pick thresholds from three different sets of values, one for each ML module inside our pipeline. Then, we compute TPR, FPR and F1-score of SLIFER with a validation set. We adopt different splitting strategies to extract the latter from our data, and we chose the thresholds that maximise the F1-score of SLIFER, and we show results in Table 6. As the table suggests, there is no general improvement by varying the size of the validation set. Thus, we take the 50% splitting strategy since we can test the found thresholds combinations with a more representative and bigger test set even though the absolute winner in terms of F1-score is the 80% strategy. The optimal thresholds vary substantially from those determined at a 1% FPR on the original datasets of the models. MalConv threshold is increased from 0.5 to 0.76, GBDT-EMBER threshold is significantly decreased from 0.82 to 0.44 and Nebula threshold is lowered from 0.97 to 0.70. Hence, we can conclude that MalConv is mostly used by SLIFER to detect malware with high-confidence, leaving samples with lower score to GBDT-EMBER (whose threshold has been halved). Thanks to this calibration, we can select the same FPR for all the models in our analysis, to produce a fair comparison. Thus, we compute ROC curves for the other models and stand-alone modules to find thresholds achieving the same FPR, as depicted in Fig. 4).
Figure 4: ROC curve of all the models under test. We take a decision threshold for each model fixing FPR at 6.7% as scored in the validation process, to have a fair comparison with SLIFER.
In the remaining part of this subsection, we use found thresholds for all models and SLIFER to compare results in a calibrated and fair experimental setting.
Take-home message 3: All the detection thresholds of SLIFER can be fine-tuned on a validation set with a simple grid-search, substantially changing the behavior of each module differently than when used standalone.
Table 6: Performance of SLIFER and optimal threshold computed on the validation set based on the best F1-Score. We report different splits of the union of $\mathcal{D}_{1}\,$and $\mathcal{D}_{2}\,$, and the correspondent best thresholds for MalConv, GBDT-EMBER, and Nebula.
Validation Split TPR FPR F1-Score $\mathcal{T}_{M}$ $\mathcal{T}_{E}$ $\mathcal{T}_{N}$
50% 0.962 0.067 0.955 0.76 0.44 0.70
60% 0.961 0.064 0.955 0.76 0.44 0.70
70% 0.960 0.067 0.953 0.76 0.50 0.70
80% 0.961 0.063 0.956 0.76 0.5 0.72
Performance on Test Set. We now analyze results on the test set, dividing metrics by dataset as we did in Sect. 5.1. Results are reported in Table 7. We compare the calibrated SLIFER with other models, setting their thresholds to the same FPR value as reported by our pipeline. We can see that on $\mathcal{D}_{1}\,$we have the same results as in the "non-calibrated" scenario, while on $\mathcal{D}_{2}\,$, our pipeline reports higher TPR and lower FPR. This results is mainly due to $\mathcal{D}_{2}\,$being more populated. Hence, the validation set contain more representative samples of each of the contained malware families. However, we can empirically state that calibrating with unbalanced datasets does not worsen results on the less present one while improving performance on the most prominent. After that, we report another effect of the calibration: on $\mathcal{D}_{2}\,$, removing dynamic analysis has a bigger impact than before, since Nebula is part of the process of fine-tuning. As regards comparisons with single models, GBDT-EMBER confirms to be the best performing one, even at this FPR, and our pipeline performs slightly worse than it in terms of TPR but with lower FPR. The error rates on the test set indicate a significant number of samples that are impossible to analyze when emulation is part of the analysis (as with Nebula, Quo.Vadis, and integral SLIFER). On the contrary, we encounter fewer errors for GBDT-EMBER and SLIFER without dynamic analysis. As ultimate test, F1-Score follows the trends already discovered earlier in this Section as well as in Table 3 using stock thresholds. Therefore, these results are consistent with our earlier findings. Calibration highlights as well that the best error management policy is to consider analysis errors as benign to ensure a usable FPR.
Results on testset families. For the sake of completeness, we report TPR on the test set for each malware family, and we report results in Table 8 and Table 9. Results are substantially aligned with the non-calibrated SLIFER. The calibration process improves detection of some of the malware families (especially the ones belonging to $\mathcal{D}_{2}\,$), like Dropper and Ransomware, while under-performing on the Rogueware family, belonging to $\mathcal{D}_{1}\,$.
Take-home message 4: The calibration of detection thresholds of SLIFER can lead to more accurate results and better robustness in general, but its impact is negligible.
Table 7: Performance of calibrated single models and SLIFER. Results in brackets are calculated considering detection errors as malware. They are considered benignware otherwise. We report the True Positive Rate (TPR), the False Positive Rate (FPR), F1 Score (F1), and the Error Rate (ER).
$\mathcal{D}_{1}\,$
Model TPR FPR F1 ER
Signatures 0.15 7.0×$\times 10^{-3}$ 0.26 0%
MalConv 0.79 2.7×$\times 10^{-2}$ 0.87 0%
GBDT 0.96 (0.97) 5.0×$\times 10^{-2}$ (0.17) 0.96 (0.93) 5.6%
Nebula 0.36 (0.70) 6.0×$\times 10^{-2}$ (0.37) 0.51 (0.71) 35%
Quo.Vadis 0.87 (0.97) 4.9×$\times 10^{-2}$ (0.29) 0.92 (0.89) 15%
SLIFER No Dyn. 0.94 (0.95) 2.3×$\times 10^{-2}$ (0.15) 0.96 (0.93) 5.5%
SLIFER 0.94 (0.98) 5.2×$\times 10^{-2}$ (0.40) 0.95 (0.87) 16%
$\mathcal{D}_{2}\,$
Model TPR FPR F1 ER
Signatures 0.35 9.6×$\times 10^{-3}$ 0.51 0%
MalConv 0.76 5.7×$\times 10^{-2}$ 0.85 0%
GBDT 0.95 (0.95) 7.1 (7.2)×$\times 10^{-2}$ 0.96 (0.96) 0.058%
Nebula 0.47 (0.92) 4.2×$\times 10^{-2}$ (0.25) 0.63 (0.89) 36%
Quo.Vadis 0.89 (0.93) 6.5×$\times 10^{-2}$ (0.14) 0.93 (0.93) 4.7%
SLIFER No Dyn. 0.89 (0.89) 4.7 (4.8)×$\times 10^{-2}$ 0.93 (0.93) 0.044%
SLIFER 0.92 (0.98) 6.1×$\times 10^{-2}$ (0.27) 0.94 (0.92) 11%
Table 8: True Positive Rate (TPR) of calibrated SLIFER divided by malware family inside in $\mathcal{D}_{1}\,$. Results in brackets are calculated considering detection errors as malware. They are considered benignware otherwise.
Architecture Backdoor Downloader Grayw. Miner Ransomw. Roguew. Spyware Virus Worm
SLIFER No Dyn. 0.98 (0.98) 0.96 (0.98) 0.93 (0.93) 1.0 (1.0) 0.99 (0.99) 0.80 (0.80) 0.96 (0.96) 0.99 (0.99) 0.99 (0.99)
SLIFER 0.98 (0.99) 0.96 (0.99) 0.94 (0.97) 1.0 (1.0) 0.99 (1.0) 0.80 (1.0) 0.96 (1.0) 0.99 (1.0) 0.99 (0.99)
Table 9: True Positive Rate (TPR) of calibrated SLIFER divided by malware family inside in $\mathcal{D}_{2}\,$. Results in brackets are calculated considering detection errors as malware. They are considered benignware otherwise.
SLIFER No Dyn. 0.99 (0.99) 0.98 (0.98) 0.92 (0.92) 0.42 (0.42) 0.97 (0.97) 0.99 (0.99) 0.98 (0.99)
SLIFER 0.99 (0.99) 0.98 (0.99) 0.92 (1.0) 0.57 (0.87) 0.99 (0.99) 0.99 (1.0) 0.99 (0.99)
5.4 Robustness Against Transfer Attacks
We now answer RQ5 by testing the robustness of our architecture against adversarial EXEmples computed on GBDT-EMBER (with its "default" decision threshold at 1% FPR) and transferred again the SLIFER, and we show results of GAMMA section-injection and padding in Table 10 and Table 11 respectively. For the sake of clarity, we split this analysis into two paragraphs.
@@ 5.4 Robustness Against Transfer Attacks
GAMMA Section Injection Attack. From the randomly-sampled 1k malicious samples, GAMMA is able to produce 388 EXEmples that bypass GBDT-EMBER detection. We use this set to compute the robustness against transfer attacks of SLIFER with and without the dynamic module. Since we are interested also in the effect that signatures play in countering malware, we want to quantify how much they can be useful when dealing with adversarial EXEmples. Thus, we also add to the comparison a version of SLIFER where YARA pattern-matching is removed (SLIFER No Sign.). Lastly, to understand whether an attack that manipulate only the structure of a program might also affect a behavioral model, we also test the robustness of Nebula alone.
Table 10: Adversarial Detection Rate (ADR) of diverse architectures/model under GAMMA Section Injection attacks. We compare not manipulated malware samples and the same samples after manipulation (Adversarial EXEmples), considering stock and calibrated architecture/models. Results in brackets are calculated considering detection errors as malware. They are considered benignware otherwise.
@@
After Calibration
Architecture DR Architecture ADR
SLIFER 1.0 (1.0) SLIFER 0.88 (0.95)
SLIFER No Dyn. 1.0 (1.0) SLIFER No Dyn. 0.81 (0.81)
SLIFER No Sign. 1.0 (1.0) SLIFER No Sign. 0.79 (0.93)
Nebula 0.55 (0.90) Nebula 0.55 (0.89)
While all untainted malware samples are detected by SLIFER, with the exception of Nebula alone due to the high number of emulation errors, results of transfer evaluation highlight an evasive trend against all models. These suggest that GAMMA attack can evade the whole SLIFER, decreasing the Detection Rate by 14%. This is achieved also thanks to the policy we established of flagging samples as benignware in case of errors of the last dynamic module. As we can see from the ADR if we count errors as malware, we derive that the 7% of Adversarial EXEmples crashes the analysis, and the remaining 7% bypasses SLIFER without pre-processing errors. By looking at Nebula performances on these data, it is likely that roughly half of them are not detected at all, but also that some are, surprisingly, evading detection. This can be noticed by looking again at the results scored by Nebula: when considering errors as malware, there is a small drop of performance (3%) implying that those EXEmples do not crash the pipeline. By inspecting the results, we discovered that (i) some malware samples obtain a different score when emulated multiple times inside a single Speakeasy session, and (ii) GAMMA changes the output score of Nebula. We analyzed those reports, and we discovered that samples that invoke specific Windows APIs like:
@@
  • GetModuleFileNameW retrieves a different file name. It is likely that Speakeasy rename the file to analyze with its hash. Thus, the output report of the untainted malware and its adversarial EXEmple counterpart are different.
This analysis is not comprehensive, as it would need deeper reverse engineering of those samples, but, even if for few samples, we can conclude that also dynamic analysis is evaded by non-behavioral attacks, not strictly due to the effect of perturbations. Lastly, 5 samples that originally caused errors during emulation are now analyzed by Speakeasy after the manipulation. This result needs further investigation, but they bring evidence that dynamic analysis based on emulation can be weak and discontinuous not only against adversarial transfer attacks, but also to multiple evaluations of the same samples.
If we remove dynamic analysis from SLIFER (SLIFER No Dyn. in Table 10), we detect a drop of ADR by 10%, implying that such a module was indeed able to stop some of the incoming attacks. However, counter-intuitively, if we instead remove signatures (SLIFER No Sign.) and we keep dynamic analysis, we assist to a drop of ADR by 15% w.r.t. SLIFER. This means that, surprisingly, pattern-matching with YARA rules has a bigger impact on robustness than dynamic analysis itself. This relevant contribution of YARA rules is due to how the section injection manipulation works: while mixing together sections extracted from input data, it is likely that GAMMA creates some patterns that trigger YARA rules, contrary to the original PE. We notice that 20 samples evade the YARA module before GAMMA manipulation, but they are detected once perturbed. To better understand this result, we analyze a small subset of rules that are triggered by adversarial EXEmples, and we analyze them through expert domain knowledge. Through this study, we isolate culprits of such triggers:
@@
Thus, we can conclude that SLIFER is indeed robust against transfer adversarial EXEmples computed against its strongest component, but such robustness is achieved more by static signatures than dynamic analysis alone. Moreover, we report another interesting effect of adversarial manipulation on signatures: some rules include a filesize condition to fire. If a PE is larger than a preset size, the rule is not fired. While this condition can be useful to avoid false alarms, it prevents some rules from being fired against adversarial EXEmples. Injected sections can increase the size of the manipulated malware samples beyond the limits set by certain rules, effectively allowing them to evade signature detection.
Calibration has a positive effect on robustness, increasing SLIFER’s ADR by 2%. As a side-effect, all the modules themselves exhibit an improved robustness, because the tuned thresholds contribute in blocking more adversarial EXEmples. In particular, the decrease of Nebula’s threshold improves by 10% its ADR, increasing its impact.
GAMMA Padding Attack. From the same 1k malware considered for GAMMA section-injection, we compute 469 adversarial EXEmples with a padding manipulation, 81 more than GAMMA sections injection attack. This difference is because padding manipulation does not modify headers of malicious samples and GBDT-EMBER can be misled more easily. Indeed, EMBER features also consider the header of the PE, and padding manipulation can fool the computation of such features. While SLIFER can detect all the original malicious samples, also when removing YARA or Nebula from its architecture, the GAMMA Padding attack decreases the Detection Rate by 13% and by 11% with stock and optimal thresholds respectively. Similarly to what we have observed for GAMMA section-injection, the calibration increases the robustness. This is caused by the decrease of the decision threshold of GBDT-EMBER, requiring attackers to create adversarial EXEmples with higher confidence. Regarding dynamic analysis, we notice the same discontinuous behaviors of Speakeasy emulation and Nebula prediction. In particular, we notice that Nebula attributes different scores to the original malware and its adversarial EXEmple counterpart. Also, two samples that originally crash emulation are correctly emulated after the attack. Another pattern that presents is the bigger contribution of YARA to ADR w.r.t. Nebula, particularly when we test stock decision thresholds. On the contrary, with calibration, Nebula has a greater impact than YARA, because its optimal threshold is lower. Contrary to sections injection attack, GAMMA padding can also succeed against YARA in two cases. We notice that two adversarial EXEmples are able to evade the first module, contrary to the original untainted malware. More precisely, the rules which fire before the padding manipulation are:
  • SUSP_XORed_URL_In_EXE: as the name tells, detects a XORed URL in an executable. The PE that triggers this rule downloads a trojan inside a system’s directory, by connecting to a XOR-obfuscated URL;
  • SUSP_RANSOMWARE_Indicator_Jul20:this rules detects the presence of specific strings like DecryptFilesHere.txt or Decrypt.txt, which are correalated with ransomware activities; [^14] https://securelist.com/lazarus-on-the-hunt-for-big-game/97757/
  • win_gandcrab_auto: this rule aims detects specific instance of the win.gandcrab malware, by looking for particular strings (automatically selected from the disassembly of memory dumps and unpacked files, using a tool named “YARA-Signator”). [^15] https://malpedia.caad.fkie.fraunhofer.de/details/win.gandcrab
Another relevant pattern is that rules triggered by adversarial EXEmples are a subset of the ones triggered by original samples. Hence, it seems that padding prevents some rules from firing. With further investigation, we discover that also the three above-mentioned rules include a condition on the filesize as described in the analysis of GAMMA section-injection. However, GAMMA padding manipulation can heavily affect PE size and consequently, they can evade signatures module as well. Hence, YARA rules can be beneficial for robustness, as seen with GAMMA sections-injection attacks, but easily bypassed by adding content if they present a limit on the file size.
Table 11: Adversarial Detection Rate (ADR) of diverse architectures/model under GAMMA Padding attacks. We compare not manipulated malware samples and the same samples after manipulation (Adversarial EXEmples), considering stock and calibrated architecture/models. Results in brackets are calculated considering detection errors as malware. They are considered benignware otherwise.
Original Malware Samples Adversarial EXEmples
No Calibration
Architecture DR Architecture ADR
SLIFER 1.0 SLIFER 0.87 (0.95)
SLIFER No Dyn. 1.0 SLIFER No Dyn. 0.80 (0.80)
SLIFER No Sign. 1.0 SLIFER No Sign. 0.77 (0.94)
Nebula 0.45 (0.81) Nebula 0.44 (0.80)
After Calibration
Architecture DR Architecture ADR
SLIFER 1.0 SLIFER 0.89 (0.95)
SLIFER No Dyn. 1.0 SLIFER No Dyn. 0.80 (0.80)
SLIFER No Sign. 1.0 SLIFER No Sign. 0.81 (0.94)
Nebula 0.52 (0.88) Nebula 0.54 (0.89)
@@
Take-home message 5: SLIFER can withstand adversarial transfer attacks, but the effect of dynamic analysis is counter-intuitively marginal. Attacks can trigger YARA rules that were not predictive against the original untainted samples, since GAMMA introduces artifacts. Also, some rules are disabled by the attack, due to constraints on filesize which are now violated by attacks.
5.5 Computation Times Comparison